# Segment anything model 2: an application to 2D and 3D medical images

### v2

@@ 1 Introduction
Medical image segmentation is a crucial task for multiple clinical applications such as disease diagnosis and clinical analysis [17, 24, 27, 33]. Despite the advancements in medical imaging technologies, segmentation remains a challenging task due to the labor-intensive nature of data annotation and the complexity of medical images [4, 21, 36].
@@
Segment Anything Model (SAM) addresses these challenges in multiple directions. On the one hand, SAM has demonstrated impressive zero-shot segmentation performance with prompt inputs, significantly reducing the need for extensive manual data annotation [6, 21]. On the other hand, several works fine-tune SAM to specific tasks and demonstrate improvements over standard segmentation techniques [10, 7, 18, 23, 19], such as nn-UNet [14]. Despite these advancements, SAM’s limitation to 2D images restricts its applicability to scenarios that require three-dimensional understanding [28].
@@
Existing work addresses this challenge by introducing extra components to SAM to enable its 3D segmentation capability. For example, SAM3D [5] combines the SAM encoder with a lightweight 3D CNN decoder; 3DSAM-A [9] modifies the original prompt encoder and mask decoder to operate in 3D; SAM-Med3D [34] introduces an additional 3D convolution before the image encoder and replaces 2D positional encoding layers with 3D one. On the other hand, the recently introduced SAM 2 [28] solves this limitation fundamentally by extending the backbone of SAM to 3D. Specifically, SAM 2 proposes a memory bank that retains information from past predictions and allows it to make predictions on slices without prompts based on the information. This feature motivates us to examine SAM 2’s ability to segment 3D medical images since video segmentation can be transferred to 3D segmentation seamlessly, i.e., each slice can be treated as a frame. Note that we will use the terms “slice” and “frame” interchangeably throughout the paper.
@@
In this paper, we extend the previous evaluation experiments on SAM [21] to SAM 2, aiming to explore the model’s effectiveness in a more complex, three-dimensional context. Specifically, we consider two evaluation pipelines: multi-frame 3D segmentation and single-frame 2D segmentation. In single-frame 2D segmentation [21], prompts are provided to the object of interest on each slice, whereas in multi-frame 3D segmentation, prompts are provided on one or a few slices selected from the volume. We further consider some unique challenges in multi-frame 3D segmentation, such as the selection of the slice(s) to be annotated, the direction of propagation, i.e., predicting slices without prompts based on ones with prompts, and the selection of prediction during propagation. Our experiments are conducted across 18 datasets, covering 5 modalities (magnetic resonance imaging (MRI), computed tomography (CT), positron emission tomography (PET), X-ray, and Ultrasound) and various body locations. We observe several trends of SAM 2 under both evaluation pipelines and summarize them in Section 5.
@@ 2.1 Evaluation of Single-Frame 2D Model Testing on 2D Images
During single-frame 2D segmentation, SAM 2 exhibits the same behavior as SAM in segmenting the object of interest based on prompts. For datasets with 2D modalities, we run SAM 2 naturally at the image level. For datasets with 3D modalities, we simulate prompts for every slice of the volume. Following previous work [21], we design single-frame 2D segmentation in a non-iterative manner in which all prompts are determined without feedback from any prior predictions. Specifically, the following four 2D prompting modes (P-Mode) are used:
@@ 2.2 Evaluation of Multi-Frame 3D Model Testing on 3D Volumes
Initial Frame Selection. In video segmentation, the initial frame to be annotated is typically the first frame of video as it arrives first in the time stream and the object of interest usually does not change shape or size dramatically between consecutive frames [35][25]. This is also the strategy used by SAM 2 for the semi-supervised video object segmentation task. However, such a strategy may be ineffective in 3D medical imaging because, in most scenarios, the boundary of the object of interest appears first, and the change between different frames can be significant. Instead, when annotating medical volumes, annotators tend to start with the slice with the object of interest being the most salient. To evaluate SAM 2’s original frame selection (choosing the first slice) and to approximate human behavior in annotation medical volumes, we consider the following three frame modes (F-Modes):
@@
Xray-Chest Montgomery County and Shenzhen Chest X-ray Datasets [15] X-ray 1 Chest 704
Xray-Hip X-ray Images of the Hip Joints [12] X-ray 2 Ilium, Femur 140
US-Breast Dataset of Breast Ultrasound Images [1] Ultrasound 1 Breast 630
US-Kidney CT2US for Kidney Segmentation [32] Ultrasound 1 Kidney 4,586
US-Muscle Transverse Musculoskeletal Ultrasound Image Segmentations [20] Ultrasound 1 Muscle 4,044
US-Ovarian-Tumor Multi-Modality Ovarian Tumor Ultrasound (MMOTU) [37] Ultrasound 1 Ovarian tumor 1,469
@@
In the 2D setting of both SAM and SAM 2, the mask predictor can create multiple mask outputs from which humans can select the one closest to their needs. If there is no human involvement in the selection step, we can simply pick the first prediction channel, which, according to previous SAM empirical results, tends to represent the smallest area among the three channels [21]. Alternatively, we can select the channel with the highest confidence based on the IoU calculated by SAM’s mask decoder.
In 3D mode, to prevent propagating ambiguity across frames, SAM 2 chooses to pick the prediction with the highest confidence and only utilizes this prediction when estimating the object of interest on other frames. However, based on previous experiments with SAM [21, 11], selecting the predictions with the highest confidence can be sub-par when compared to selecting the first channel in the medical imaging field. This observation motivates us with the following predicted mask selection strategies:
@@
MRI-Spine Spinal Cord Grey Matter Segmentation Challenge [26] MRI 2 Gray matter and spinal cord 551 40
MRI-Heart Medical Segmentation Decathlon [31] MRI 1 Heart 1,301 20
MRI-Prostate Initiative for Collaborative Computer Vision Benchmarking [16] MRI 1 Prostate 1854 115
MRI-Brain The Multimodal Brain Tumor Image Segmentation Benchmark (BraTS) [22] MRI 3 GD-enhancing tumor, Peritumoral edema, necrotic and non-enhancing tumor core 12,591 206
MRI-Breast Duke Breast Cancer MRI: Breast + FGT Segmentation[30, 13] MRI 2 Vessel and fibroglandular tissue 14438 100
CT-Colon Medical Segmentation Decathlon [31] CT 1 Colon cancer primaries 1,285 126
CT-HepaticVessel Medical Segmentation Decathlon [31] CT 1 Vessels 13,046 303
CT-Pancreas Medical Segmentation Decathlon [31] CT 1 Parenchyma and mass 8,792 281
CT-Spleen Medical Segmentation Decathlon [31]) CT 1 Spleen 1,051 41
CT-Liver The Liver Tumor Segmentation Benchmark (LiTS) [3] CT 1 Liver 5,501 131
CT-Organ CT Volumes with Multiple Organ Segmentations (CT-ORG) [29] CT 5 Liver, bladder, lungs, kidney, and bone 4,776 10
PET-Whole-Body A FDG-PET/CT dataset with annotated tumor lesions [8] PET/CT 1 Lesion 1,015 42
@@ 3 Dataset
Consistent with the previous experimental study on SAM [21], this study utilizes 18 diverse medical datasets to evaluate the performance of SAM 2. Specifically, all datasets are evaluated during single-frame 2D segmentation, and twelve datasets with 3D modalities (MRI, CT, and PET) are evaluated during multi-frame 3D segmentation. The “US-Nerve” dataset[2], which was included in the previous SAM evaluation paper[21], has been removed from this project due to licensing regulations. We also updated the MRI-Breast dataset to include more annotated volumes. The pre-processing steps are consistent with the previous study, except that for SAM 2 3D mode, we have to convert the input images to JPG format instead of PNG. We also provided visual representations of the annotations masks for each dataset in Fig. 2.

### v3

@@ 1 Introduction
Medical image segmentation is a crucial task for multiple clinical applications such as disease diagnosis and clinical analysis [lew2024publicly, patil2013medical, ramesh2021review, wang2024sam]. Despite the advancements in medical imaging technologies, segmentation remains a challenging task due to the labor-intensive nature of data annotation and the complexity of medical images [bran2024qubiq, mazurowski2023segment, yin2018medical].
@@
Segment Anything Model (SAM) addresses these challenges in multiple directions. On the one hand, SAM has demonstrated impressive zero-shot segmentation performance with prompt inputs, significantly reducing the need for extensive manual data annotation [cheng2023sam, mazurowski2023segment]. On the other hand, several works fine-tune SAM to specific tasks and demonstrate improvements over standard segmentation techniques [gu2024segmentanybone, fu2023effectiveness, li2024polyp, na2024segment, ma2024segment], such as nn-UNet [isensee2021nnu]. Despite these advancements, SAM’s limitation to 2D images restricts its applicability to scenarios that require three-dimensional understanding [ravi2024sam2].
@@
Existing work addresses this challenge by introducing extra components to SAM to enable its 3D segmentation capability. For example, SAM3D [bui2023sam3d] combines the SAM encoder with a lightweight 3D CNN decoder; 3DSAM-A [gong20233dsam] modifies the original prompt encoder and mask decoder to operate in 3D; SAM-Med3D [wang2023sammed3d] introduces an additional 3D convolution before the image encoder and replaces 2D positional encoding layers with 3D one. On the other hand, the recently introduced SAM 2 [ravi2024sam2] solves this limitation fundamentally by extending the backbone of SAM to 3D. Specifically, SAM 2 proposes a memory bank that retains information from past predictions and allows it to make predictions on slices without prompts based on the information. This feature motivates us to examine SAM 2’s ability to segment 3D medical images since video segmentation can be transferred to 3D segmentation seamlessly, i.e., each slice can be treated as a frame. Note that we will use the terms “slice” and “frame” interchangeably throughout the paper.
@@
In this paper, we extend the previous evaluation experiments on SAM [mazurowski2023segment] to SAM 2, aiming to explore the model’s effectiveness in a more complex, three-dimensional context. Specifically, we consider two evaluation pipelines: multi-frame 3D segmentation and single-frame 2D segmentation. In single-frame 2D segmentation [mazurowski2023segment], prompts are provided to the object of interest on each slice, whereas in multi-frame 3D segmentation, prompts are provided on one or a few slices selected from the volume. We further consider some unique challenges in multi-frame 3D segmentation, such as the selection of the slice(s) to be annotated, the direction of propagation, i.e., predicting slices without prompts based on ones with prompts, and the selection of prediction during propagation. Our experiments are conducted across 18 datasets, covering 5 modalities (magnetic resonance imaging (MRI), computed tomography (CT), positron emission tomography (PET), X-ray, and Ultrasound) and various body locations. We observe several trends of SAM 2 under both evaluation pipelines and summarize them in Section 5.
@@ 2.1 Evaluation of Single-Frame 2D Model Testing on 2D Images
During single-frame 2D segmentation, SAM 2 exhibits the same behavior as SAM in segmenting the object of interest based on prompts. For datasets with 2D modalities, we run SAM 2 naturally at the image level. For datasets with 3D modalities, we simulate prompts for every slice of the volume. Following previous work [mazurowski2023segment], we design single-frame 2D segmentation in a non-iterative manner in which all prompts are determined without feedback from any prior predictions. Specifically, the following four 2D prompting modes (P-Mode) are used:
@@ 2.2 Evaluation of Multi-Frame 3D Model Testing on 3D Volumes
Initial Frame Selection. In video segmentation, the initial frame to be annotated is typically the first frame of video as it arrives first in the time stream and the object of interest usually does not change shape or size dramatically between consecutive frames [wang2021unidentified][pont20172017]. This is also the strategy used by SAM 2 for the semi-supervised video object segmentation task. However, such a strategy may be ineffective in 3D medical imaging because, in most scenarios, the boundary of the object of interest appears first, and the change between different frames can be significant. Instead, when annotating medical volumes, annotators tend to start with the slice with the object of interest being the most salient. To evaluate SAM 2’s original frame selection (choosing the first slice) and to approximate human behavior in annotation medical volumes, we consider the following three frame modes (F-Modes):
@@
Xray-Chest Montgomery County and Shenzhen Chest X-ray Datasets [jaeger2014two] X-ray 1 Chest 704
Xray-Hip X-ray Images of the Hip Joints [gut2021x] X-ray 2 Ilium, Femur 140
US-Breast Dataset of Breast Ultrasound Images [al2020dataset] Ultrasound 1 Breast 630
US-Kidney CT2US for Kidney Segmentation [song2022ct2us] Ultrasound 1 Kidney 4,586
US-Muscle Transverse Musculoskeletal Ultrasound Image Segmentations [marzola2021deep] Ultrasound 1 Muscle 4,044
US-Ovarian-Tumor Multi-Modality Ovarian Tumor Ultrasound (MMOTU) [zhao2022multi] Ultrasound 1 Ovarian tumor 1,469
@@
In the 2D setting of both SAM and SAM 2, the mask predictor can create multiple mask outputs from which humans can select the one closest to their needs. If there is no human involvement in the selection step, we can simply pick the first prediction channel, which, according to previous SAM empirical results, tends to represent the smallest area among the three channels [mazurowski2023segment]. Alternatively, we can select the channel with the highest confidence based on the IoU calculated by SAM’s mask decoder.
In 3D mode, to prevent propagating ambiguity across frames, SAM 2 chooses to pick the prediction with the highest confidence and only utilizes this prediction when estimating the object of interest on other frames. However, based on previous experiments with SAM [mazurowski2023segment, gu2024build], selecting the predictions with the highest confidence can be sub-par when compared to selecting the first channel in the medical imaging field. This observation motivates us with the following predicted mask selection strategies:
@@
MRI-Spine Spinal Cord Grey Matter Segmentation Challenge [prados2017spinal] MRI 2 Gray matter and spinal cord 551 40
MRI-Heart Medical Segmentation Decathlon [simpson2019large] MRI 1 Heart 1,301 20
MRI-Prostate Initiative for Collaborative Computer Vision Benchmarking [lemaitre2015computer] MRI 1 Prostate 1854 115
MRI-Brain The Multimodal Brain Tumor Image Segmentation Benchmark (BraTS) [menze2014multimodal] MRI 3 GD-enhancing tumor, Peritumoral edema, necrotic and non-enhancing tumor core 12,591 206
MRI-Breast Duke Breast Cancer MRI: Breast + FGT Segmentation[saha2018machine, hu2022fully] MRI 2 Vessel and fibroglandular tissue 14438 100
CT-Colon Medical Segmentation Decathlon [simpson2019large] CT 1 Colon cancer primaries 1,285 126
CT-HepaticVessel Medical Segmentation Decathlon [simpson2019large] CT 1 Vessels 13,046 303
CT-Pancreas Medical Segmentation Decathlon [simpson2019large] CT 1 Parenchyma and mass 8,792 281
CT-Spleen Medical Segmentation Decathlon [simpson2019large]) CT 1 Spleen 1,051 41
CT-Liver The Liver Tumor Segmentation Benchmark (LiTS) [bilic2023liver] CT 1 Liver 5,501 131
CT-Organ CT Volumes with Multiple Organ Segmentations (CT-ORG) [rister2019ct] CT 5 Liver, bladder, lungs, kidney, and bone 4,776 10
PET-Whole-Body A FDG-PET/CT dataset with annotated tumor lesions [gatidis2022whole] PET/CT 1 Lesion 1,015 42
@@ 3 Dataset
Consistent with the previous experimental study on SAM [mazurowski2023segment], this study utilizes 18 diverse medical datasets to evaluate the performance of SAM 2. Specifically, all datasets are evaluated during single-frame 2D segmentation, and twelve datasets with 3D modalities (MRI, CT, and PET) are evaluated during multi-frame 3D segmentation. The “US-Nerve” dataset[anna2016ultrasound], which was included in the previous SAM evaluation paper[mazurowski2023segment], has been removed from this project due to licensing regulations. We also updated the MRI-Breast dataset to include more annotated volumes. The pre-processing steps are consistent with the previous study, except that for SAM 2 3D mode, we have to convert the input images to JPG format instead of PNG. We also provided visual representations of the annotations masks for each dataset in Fig. 2.
@@
3.3 Video Dataset
yaqian[update video dataset]
Table 3: Video datasets evaluated in this paper: “num. masks” refers to the number of images with non-zero masks.
@@ 4.2.5 Comparison between Single-frame and Multi-frame Evaluation
4.3 Comparison between SAM2 and Other 3D segmentation
4.3.1 Comparison between SAM2 and SAM-Med3D
4.3.2 Comparison between SAM2 and Cuite
4.4 Interactive Mode